# DeepAndes: A Self-Supervised Vision Foundation Model for Multi-Spectral Remote Sensing Imagery of the Andes

### v1

@@ Abstract
By mapping sites at large scales using remotely sensed data, archaeologists can generate unique insights into long-term demographic trends, inter-regional social networks, and past adaptations to climate change. Remote sensing surveys complement field-based approaches, and their reach can be especially great when combined with deep learning and computer vision techniques. However, conventional supervised deep learning methods face challenges in annotating fine-grained archaeological features at scale. While recent vision foundation models have shown remarkable success in learning large-scale remote sensing data with minimal annotations, most off-the-shelf solutions are designed for RGB images rather than multi-spectral satellite imagery, such as the 8-band data used in our study. In this paper, we introduce DeepAndes, a transformer-based vision foundation model trained on three million multi-spectral satellite images, specifically tailored for Andean archaeology. DeepAndes incorporates a customized DINOv2 self-supervised learning algorithm optimized for 8-band multi-spectral imagery, marking the first foundation model designed explicitly for the Andes region. We evaluate its image understanding performance through imbalanced image classification, image instance retrieval, and pixel-level semantic segmentation tasks. Our experiments show that DeepAndes achieves superior F1 scores, mean average precision, and Dice scores in few-shot learning scenarios, significantly outperforming models trained from scratch or pre-trained on smaller datasets. This underscores the effectiveness of large-scale self-supervised pre-training in archaeological remote sensing. Codes will be available on https://github.com/geopacha/DeepAndes.
Index Terms:
Foundation Model, Self-supervised Learning, DINOv2, Multi-Spectral Imaging, Remote Sensing, Andean Archaeology
@@ I Introduction
With the rapid advancement of artificial intelligence, deep learning has made significant contributions in a diverse series of domains [29, 30, 50, 40, 49, 53]. Recently, the emergence of “foundation models” (FMs) has further expanded the scope of deep learning applications [51, 13, 23, 22, 4, 12], including remote sensing [32, 33, 31]. As summarized in [33], foundation models pretrained on massive remote sensing datasets demonstrate strong adaptability across a wide range of downstream tasks in both earth and social sciences. As a domain-specific application, archaeological remote sensing has particularly benefited from foundation models, which have shown promise in tasks such as artifact recognition [5], detection of archaeological structures [19], and texture analysis [1], offering new opportunities for large-scale archaeological analysis. Despite these advances, most of the earth’s surface has yet to be systematically surveyed, and vast regions are underrepresented in archaeological research, including the Andes region. High levels of inter-regional variability in land cover and the diversity of archaeological features themselves have proven to be major barriers to achieving such coverage. To the best of our knowledge, no foundation model has been specifically developed for Andean archaeology using 8-channel hyperspectral satellite imagery. Thus, developing a new AI foundation model with broad utility across the social and earth sciences in the Andes would be highly valuable, enabling experts to contribute where they are most effective—as observers and analysts in the archaeological workflow.
@@
To bridge this gap, this paper proposes DeepAndes, the first vision foundation model for remote sensing of the Andean region. As shown in Fig. 1, the multi-spectral training dataset, sourced from the central Andes (Figs. 1a-c), covers 488,640 km2 and encompasses 8 distinct land cover types (Fig. 1c). DeepAndes, a 307M-parameter vision transformer (ViT) model, is trained on 3 million satellite image patches. A key challenge in developing a novel vision foundation model learning framework for high-dimensional multi-spectral remote sensing images is that most off-the-shelf vision foundation models are primarily designed for natural images with only RGB bands (red, green, and blue). To address this problem, we customize the DINOv2 [36] self-supervised learning (SSL) pipelines to support our 8-band multi-spectral imagery. During pre-training, we also adjusted the scale of global-local image views to align with the size of architectural and archaeological features in the dataset. As shown in Fig. 1e, to evaluate DeepAndes, we investigated its image pattern recognition and few-shot learning capabilities through three key vision-based remote sensing tasks: imbalanced image classification, image instance retrieval, and image segmentation. These downstream tasks are crucial for large-scale remote sensing datasets, where data are often highly heterogeneous and features are sparsely distributed. Similarly to [39], DeepAndes-based image instance retrieval can help identify patterns in the distribution of cultural and natural features within vast datasets, improving the efficiency of large-scale archaeological analysis.
@@
Through extensive experiments employing archaeological features as a test case, our results demonstrate that self-supervised pre-training enhances DeepAndes’s performance in both image-level and pixel-level downstream tasks. Among the imbalanced archaeological loci (discrete archaeological features of interest) classification task, DeepAndes achieves an F1 score of 0.886. Even when fine-tuned on only 10% of the original classification data, it maintains a strong F1 score of 0.83, substantially outperforming models trained from scratch using supervised methods. For the archaeological image instance retrieval task, features extracted from the frozen DeepAndes backbone achieve a mean average precision (mAP) of 0.869 within the top-k retrieved samples (mAP@$k$, k=5), demonstrating the effectiveness of the feature representations. For pixel-level dense feature recognition, we evaluated three challenging archaeological loci segmentation tasks, where the diversity of locus objects adds complexity. Our few-shot learning results show that DeepAndes excels in both transfer learning and fine-tuning settings when using a simple linear segmentation head. Interestingly, the scaling law is observed from the DeepAndes model as depicted in Fig. 2. In short, DeepAndes, pre-trained on 3 million diverse remote sensing satellite image patches, outperforms the same ViT models pre-trained on smaller-scale datasets (e.g., 30K and 300K images) across all downstream tasks, underscoring the effectiveness of large-scale self-supervised learning in this work.
@@
  • We revise the DINOv2 framework for multi-spectral pre-training using 8-band WorldView-2 and Worldview-3 satellite imagery. This modified framework leverages the flexibility of DINOv2 with data preprocessing, data augmentation, and network architecture, enabling seamless adaptation to other multi-channel remote sensing data in the future, regardless of the number of input channels.
@@ II-B Self-Supervised Representation Learning Strategies
In the field of remote sensing and computer vision, self-supervised learning represents a fundamental component in foundation model pre-training [33]. During the pre-training stage, SSL allows models to learn meaningful representations without relying on labeled datasets. This capability is particularly beneficial in remote sensing, where labeled data is often scarce. Models pre-trained using SSL are adept at identifying patterns and features in large volumes of unlabeled remote sensing data, making them highly effective for various downstream applications. As summarized in [33], some commonly used SSL techniques in remote sensing FMs can be categorized into two main types: contrastive learning and predictive coding.
Contrastive Learning. Contrastive learning method focuses on learning data representations by comparing different augmented versions of the same data point. It pulls similar (positive) pairs closer and push dissimilar (negative) pairs apart in the representation space. The methodology depends on advanced data augmentation techniques, such as random cropping, rotation, and color jittering, to create diverse views of the same image. Notable implementations that have shown success in remote sensing applications include DINO [8], which employs self-distillation, SimCLR [10], which uses a simple contrastive framework, and MOCO (Momentum Contrast) [25], which leverages a dynamic dictionary approach [33].
@@ III-A Million-Scale Training Dataset Construction
As described previously in Figure 1, the dataset used for SSL pre-training was sourced from the entire Peruvian Andes, surveyed from 6 teams, spanning 488,640 $km^{2}$, and including 8 distinct land cover types. Our high-resolution satellite imagery is from the WorldView-2 (WV-2) and WorldView-3 (WV-3) satellites, both of which are composed of eight multi-spectral bands, plus a panchromatic band. After pansharpening (fusing the panchromatic band with other spectral bands), WV-2 imagery is of a Ground Sample Distance (GSD) of 0.46 m (at nadir), while WV-3 imagery provides 0.31 m (at nadir) GSD resolution. Our imagery processing pipeline includes steps for image orthorectification, cloud removal, and GSD commensuration. Image processing includes upsampling WV-2 imagery to match the GSD of WV-3 imagery. To construct a diverse, large-scale pre-training dataset, we randomly sampled the Central Andean area (encompassing western Peru and parts of northern Chile and northwestern Bolivia) into non-overlapping 256 $\times$ 256 image patches (shown as the red box in Fig. 1). Featureless images with entirely dark or bright pixels, such as those containing only water or clouds, were removed. This process yielded a total of 3 million 256 $\times$ 256 image patches for pre-training DeepAndes.
@@ III-B2 SSL Pre-training
In our work, we selected the ViT “large” (ViT-L/14) with 307M parameters as the backbone. The default multi-crop DINOv2 training uses a scale factor $s=0.32$ to define the scaling range of local crop as (0.05, $s$), and ($s$, 1) for global views. In our study, we applied the scaling range (0.2, 0.5) for the local view and (0.5, 1) for the global view, due to the sparse image feature distribution and noise present in our remote sensing images. The other default hyperparameters for DINOv2 training algorithm were used for DeepAndes, as detailed in [36] with the following modifications: a base learning rate of 0.0002, and the entire training process covering approximately 40 times passes over the 3-million-image dataset. DeepAndes was trained using AdamW ($\beta_{1}=0.9,\beta_{2}=0.999$) with float16 precision. For ViT-L, we used 65,536 prototypes, resulting in 65,536-dimensional projection heads. Distributed data parallel (DDP) training was employed across 8 NVIDIA DGX-A100 nodes, with a batch size of 64 per GPU node. To implement complex augmentation pipelines, we adapted the default augmentation modules for 8-band image data with minor modifications.
@@ III-B3 Feature Embeddings
For an input image of size 256 $\times$ 256, it is resized to 224 $\times$ 224 before being input to the model, in accordance with the default settings of DINOv2. Using ViT-L with a patch size of 14 $\times$ 14, the pre-trained DeepAndes projects the input image into vector representations (tokens) of dimension 1,024 in the feature space. With the default settings, this includes one class token and 256 image patch tokens. These vectors generated by foundation models present meaningful feature information of the input images with a reduced dimension can be further utilized for diverse downstream tasks [44, 45, 24]. In our case, the class token can be used for downstream tasks such as image classification by concatenating a simple linear classifier. The patch tokens can be utilized for pixel-level image recognition tasks, such as semantic segmentation, by connecting them to a segmentation head.
@@ III-C1 Handling Imbalance in Archaeological Data
Another major challenge in archaeological remote sensing is the inherent imbalance in archaeological data, particularly in our remote sensing satellite imagery. One of the previous studies [49] shows that the proportion of our satellite imagery that include archaeological settlement features is typically low, at less than 7% in the surveyed areas. For example, in a manually labeled dataset consisting of 5,830 labeled images sourced from a 4,000 km2 survey area, the ratio of “positive” (containing settlement features) to “negative” (not containing settlement features) images can be as low as 1:100. The pre-trained model can be adapted and fine-tuned to improve performance in minority classes when the downstream task is bottlenecked by data imbalance and scarcity.
@@ III-C2 Archaeological Downstream Tasks
To support the motivation above, we evaluate DeepAndes on three image understanding downstream tasks in archaeology: image-level classification and retrieval, and pixel-level segmentation, as shown in the Figure 4.
Imbalanced Loci Classification. As described in the previous section, archaeological loci, such as settlement buildings and corrals, are sparsely distributed across the Andes area. Thus, we determine the first image-level downstream task as imbalanced loci classification. Specifically, images of the classification dataset are categorized into two classes: “positive” (i.e., the presence of loci) and “negative” (i.e., no presence of loci), denoted as background images. The class token of the pre-trained ViT-L backbone was connected to a simpler linear classifier with two fully connected layers. Given an input image, the fine-tuned classifier predicts either “positive” or “negative.” To represent the data imbalance in this task, the ratio of “positive” to “negative” images was set to 1:10. Additionally, to evaluate the few-shot learning capability of the pre-trained transformer backbones, we assess the classification task at different downstream dataset scales (10%, 30%, 50%, and 100% of the original classification dataset).
Image Instance Retrieval. Another image-level downstream task we perform is the image instance retrieval task. Unlike classification, image retrieval does not require fine-tuning on DeepAndes. Instead, we use an archaeological loci image as the input query and retrieve the top-k most similar images from the database, ranking them based on their cosine similarity to the query loci image (via the class token). The performance of the model is then evaluated by the mean average precision of the matched image class (“positive” or “negative”) in the retrieved results.
Few-shot Loci Segmentation In this work, we also consider the task of recognizing dense features, such as semantic segmentation at the pixel level. Specifically, we perform few-shot semantic segmentation tasks for three types of loci: active buildings, active corrals, and archaeological corrals. Patch-level features are extracted from the images via patch tokens and then concatenated with a simple linear segmentation head. Unlike the loci classification task, pixel-level semantic segmentation enables more precise localization of archaeological features within the image. As described in [36], we freeze the pre-trained transformer backbone and only fine-tune the linear segmentation head to generate output logits from the patch tokens.

### v2

@@ Abstract
By mapping sites at large scales using remotely sensed data, archaeologists can generate unique insights into long-term demographic trends, inter-regional social networks, and human adaptations in the past. Remote sensing surveys complement field-based approaches, and their reach can be especially great when combined with deep learning and computer vision techniques. However, conventional supervised deep learning methods face challenges in annotating fine-grained archaeological features at scale. In addition, while recent vision foundation models have shown remarkable success in learning large-scale remote sensing data with minimal annotations, most off-the-shelf solutions are designed for RGB images rather than multi-spectral satellite imagery, such as the 8-band data used in our study. In this paper, we introduce DeepAndes, a transformer-based vision foundation model trained on three million multi-spectral satellite images, specifically tailored for Andean archaeology. DeepAndes incorporates a customized DINOv2 self-supervised learning algorithm optimized for 8-band multi-spectral imagery, marking the first foundation model designed explicitly for the Andes region. We evaluate its image understanding performance through imbalanced image classification, image instance retrieval, and pixel-level semantic segmentation tasks. Our experiments show that DeepAndes achieves superior F1 scores, mean average precision, and Dice scores in few-shot learning scenarios, significantly outperforming models trained from scratch or pre-trained on smaller datasets. This underscores the effectiveness of large-scale self-supervised pre-training in archaeological remote sensing. Codes will be available on https://github.com/geopacha/DeepAndes.
@@ I Introduction
With the rapid advancement of artificial intelligence, deep learning has made significant contributions in a diverse series of domains [30, 32, 53, 42, 52, 56]. Recently, the emergence of foundation models (FMs) has further expanded the scope of deep learning applications [54, 13, 23, 22, 4, 12], including remote sensing [34, 35, 33]. As summarized in [35], foundation models pretrained on massive remote sensing datasets demonstrate strong adaptability across a wide range of downstream tasks in both earth and social sciences. As a domain-specific application, archaeological remote sensing has particularly benefited from foundation models, which have shown promise in tasks such as artifact recognition [5], detection of archaeological structures [19], and texture analysis [1], offering new opportunities for large-scale archaeological analysis. Despite these advances, most of the earth’s surface has yet to be systematically surveyed, and vast regions are underrepresented in archaeological research, including the Andes region. High levels of inter-regional variability in land cover and the diversity of archaeological features themselves have proven to be major barriers to achieving such coverage. Thus, developing a new AI foundation model with broad utility across the social and earth sciences in the Andes would be highly valuable, enabling experts to contribute where they are most effective—as observers and analysts in the archaeological workflow.
@@
To bridge this gap, this paper proposes DeepAndes, the first vision foundation model for remote sensing of the Andean region. To the best of our knowledge, no foundation model has been specifically developed for the Andes using 8-channel multi-spectral satellite imagery. As shown in Fig. 1, the multi-spectral training dataset, sourced from the central Andes (Figs. 1a-c), covers 488,640 km2 and encompasses 8 distinct land cover types (Fig. 1c). DeepAndes, a 307M-parameter vision transformer (ViT) model, is trained on 3 million satellite image patches. A key challenge in developing a novel vision foundation model learning framework for high-dimensional multi-spectral remote sensing images is that most off-the-shelf vision foundation models are primarily designed for three-band (RGB) imagery. To address this problem, we customized the DINOv2 [38] self-supervised learning (SSL) pipelines to support our 8-band multi-spectral imagery. During pre-training, we also adjusted the scale of global-local image views to align with the size of architectural and archaeological features in the dataset. As shown in Fig. 1e, to evaluate DeepAndes, we investigated its image pattern recognition and few-shot learning capabilities through three key vision-based remote sensing tasks: imbalanced image classification, image instance retrieval, and image segmentation. These downstream tasks are crucial for large-scale remote sensing datasets, where data are often highly heterogeneous and features are sparsely distributed. Similarly to [41], DeepAndes-based image instance retrieval can help identify patterns in the distribution of cultural and natural features within vast datasets, improving the efficiency of large-scale archaeological analysis.
@@
Through extensive experiments employing archaeological features as a test case, our results demonstrate that self-supervised pre-training enhances DeepAndes’s performance in both image-level and pixel-level downstream tasks. Among the imbalanced archaeological loci (discrete archaeological features of interest) classification task, DeepAndes achieves an F1 score of 0.886. Even when fine-tuned on only 10% of the original classification data, it maintains a strong F1 score of 0.83, substantially outperforming models trained from scratch using supervised methods. For the archaeological image instance retrieval task, features extracted from the frozen DeepAndes backbone achieve a mean average precision (mAP) of 0.869 within the top-k retrieved samples (mAP@$k$, k=5), demonstrating the effectiveness of the feature representations. For pixel-level dense feature recognition, we evaluated segmentation tasks across three challenging archaeological loci, where the diversity of locus (singular form of loci) objects adds complexity. Our few-shot learning results show that DeepAndes excels in both transfer learning and fine-tuning settings when using a simple linear segmentation head. Interestingly, the scaling law is observed from the DeepAndes model as depicted in Fig. 2. In short, DeepAndes, pre-trained on 3 million diverse remote sensing satellite image patches, outperforms the same ViT models pre-trained on smaller-scale datasets (e.g., 30K and 300K images) across all downstream tasks, underscoring the effectiveness of large-scale self-supervised learning in this work.
@@
  • We revise the DINOv2 framework for multi-spectral pre-training using 8-band WorldView-2 (WV-2) and Worldview-3 (WV-3) satellite imagery. This modified framework leverages the flexibility of DINOv2 with data preprocessing, data augmentation, and network architecture, enabling seamless adaptation to other multi-channel remote sensing data in the future, regardless of the number of input channels.
@@ II-B Self-Supervised Representation Learning Strategies
In the field of remote sensing and computer vision, self-supervised learning represents a fundamental component in foundation model pre-training [35]. During the pre-training stage, SSL allows models to learn meaningful representations without relying on labeled datasets. This capability is particularly beneficial in remote sensing, where labeled data is often scarce. Models pre-trained using SSL are adept at identifying patterns and features in large volumes of unlabeled remote sensing data, making them highly effective for various downstream applications. As summarized in [35], commonly used SSL techniques in remote sensing FMs can fall into two main categories: contrastive learning and predictive coding.
Contrastive Learning. Contrastive learning focuses on learning data representations by comparing different augmented versions of the same data point. It pulls similar (positive) pairs closer and push dissimilar (negative) pairs apart in the representation space. The methodology depends on advanced data augmentation techniques, such as random cropping, rotation, and color jittering, to create diverse views of the same image. Notable implementations that have shown success in remote sensing applications include DINO [8], which employs self-distillation, SimCLR [10], which uses a simple contrastive framework, and MOCO (Momentum Contrast) [25], which leverages a dynamic dictionary approach [35].
@@ III-A Million-Scale Training Dataset Construction
As described previously in Figure 1, the dataset used for SSL pre-training was sourced from across the Peruvian Andes, surveyed from 6 teams, spanning 488,640 $km^{2}$, and including 8 distinct land cover types. Our high-resolution satellite imagery is from the WorldView-2 and WorldView-3 satellites, both of which are composed of eight multi-spectral bands, plus a panchromatic band. After pansharpening (fusing the panchromatic band with other spectral bands), WV-2 imagery is of a Ground Sample Distance (GSD) of 0.46 m (at nadir), while WV-3 imagery provides 0.31 m (at nadir) GSD resolution. Our imagery processing pipeline includes steps for image orthorectification, cloud removal, and GSD commensuration. Image processing includes upsampling WV-2 imagery to match the GSD of WV-3 imagery. To construct a diverse, large-scale pre-training dataset, we randomly sampled the Central Andean area (encompassing western Peru and parts of northern Chile and northwestern Bolivia) into non-overlapping 256 256 image patches (shown as the red box in Fig. 1). Featureless images with entirely dark or bright pixels, such as those containing only water or clouds, were removed. This process yielded a total of 3 million 256 256 image patches for pre-training DeepAndes.
@@ III-B2 SSL Pre-training
In our work, we selected the ViT “large” (ViT-L/14) with 307M parameters as the backbone. The default multi-crop DINOv2 training uses a scale factor $s=0.32$ to define the scaling range of local crop as (0.05, $s$), and ($s$, 1) for global views. In our study, we applied the scaling range (0.2, 0.5) for the local view and (0.5, 1) for the global view, due to the sparse image feature distribution and noise present in our remote sensing images. The other default hyperparameters for DINOv2 training algorithm were used for DeepAndes, as detailed in [38] with the following modifications: a base learning rate of 0.0002, and the entire training process covering approximately 40 passes over the 3-million-image dataset. DeepAndes was trained using AdamW ($\beta_{1}=0.9,\beta_{2}=0.999$) with float16 precision. For ViT-L, we used 65,536 prototypes, resulting in 65,536-dimensional projection heads. Distributed data parallel (DDP) training was employed across 8 NVIDIA DGX-A100 nodes, with a batch size of 64 per GPU node. To implement complex augmentation pipelines, we adapted the default augmentation modules for 8-band image data with minor modifications. The summary of hyperparameters for pre-training 3 million image crops is provided in Table I.
@@ III-B3 Feature Embeddings
For an input image of size 256 256, it is resized to 224 224 before being input to the model, in accordance with the default settings of DINOv2. Using ViT-L with a patch size of 14 14, the pre-trained DeepAndes projects the input image into vector representations (tokens) of dimension 1,024 in the feature space. With the default settings, this includes one class token and 256 image patch tokens. These vectors generated by foundation models present meaningful feature information of the input images with a reduced dimension can be further utilized for diverse downstream tasks [47, 48, 24]. In our case, the class token can be used for downstream tasks such as image classification by concatenating a simple linear classifier. The patch tokens can be utilized for pixel-level image recognition tasks, such as semantic segmentation, by connecting them to a segmentation head.
TABLE I: Summary of pre-training hyperparameters. Other settings are default as [38]
Category Key Settings Architecture ViT-Large; Input channels = 8; Patch size = 14; 0.3B parameters. Student Drop path rate = 0.3; Layerscale = 1e-5; Drop path uniform = true; FFN layer = MLP; qkvqkv, projection, and FFN bias = true; No register tokens; Interpolate offset = 0.1. Teacher Momentum = 0.992 1.0; temperature = 0.04 0.07; Warmup epochs = 30. Training Setup Epochs = 500; Iterations per epoch = 1000; Batch size per GPU = 64; GPUs = 8; Warmup epochs = 80; Gradient clip = 3.0. Optimization Base LR = 2e-4 (scaled $\sqrt{1024/\text{batch}}$); Weight decay = 0.04 0.4; Layerwise decay = 0.9; AdamW $(\beta_{1},\beta_{2})=(0.9,0.999)$. Loss / Heads DINO and iBOT loss weights = 1.0 each; iBOT mask ratio = 0.1–0.5; Separate iBOT head = true; Prototypes = 65,536; Head bottleneck dim = 256; Hidden dim = 2,048; Layers = 3. Global-Local Crops Global crops: size = 224, scale = 0.5–1.0; Local crops: size = 98 (divisible by 14), number = 8, scale = 0.2–0.5. Mixed precision fp16 for parameters and reduce; fp32 for buffers
@@ III-C1 Handling Imbalance in Archaeological Data
Another major challenge in archaeological remote sensing is the inherent imbalance of archaeological data. One of the previous studies [52] shows that the proportion of our satellite imagery that include archaeological settlement features is typically low, at less than 7% in the surveyed areas. For example, in a manually labeled dataset consisting of 5,830 labeled images sourced from a 4,000 km2 survey area, the ratio of “positive” (containing settlement features) to “negative” (not containing settlement features) images can be as low as 1:100.
To address the data imbalance, common approaches include data resampling or ensemble methods [31], typically rely on extensive domain expertise or labeled data and are difficult to scale effectively. Foundation models leverage large-scale unlabeled data through task-agnostic pre-training, which has been shown in [27, 43] to alleviate data imbalance and reduce overfitting in downstream tasks. K-fold cross-validation was further employed to validate the robustness of this approach in addressing the imbalance scenario.
@@ III-C2 Archaeological Downstream Tasks
To support the motivation above, we evaluate DeepAndes on three image understanding downstream tasks in archaeology: image-level classification and retrieval, and pixel-level segmentation, as shown in Figure 4.
Imbalanced Loci Classification. Loci classification, or archaeological data interest classification, is the process of distinguishing image crops with high archaeological relevance from those that are less informative. This step is crucial in large-scale remote sensing surveys, where vast amounts of satellite imagery include diverse content—natural landforms, vegetation, and modern infrastructure—of which only a small fraction is archaeologically significant. Effective classification of Andean corrals and buildings directs surveyors to the most relevant imagery, saving their time and improving archaeological interpretation. Specifically, images in the classification dataset were labeled as either “positive” (loci present) or “negative” (no presence of loci). The class token of the pre-trained ViT-L backbone was connected to a linear classifier with two fully connected layers to predict these classes. To reflect data imbalance, the positive-to-negative ratio was set to 1:10. Few-shot performance was then evaluated at different dataset scales (10%, 30%, 50%, and 100%).
Image Instance Retrieval. Another image-level downstream task conducted in this study is image instance retrieval. Unlike classification, image retrieval does not require fine-tuning on DeepAndes. Instead, we use an image of archaeological loci as the input query and retrieve the top-k most similar images from the database, ranking them based on their cosine similarity to the query loci image (via the class token). Image retrieval enables rapid construction of relevant datasets without additional supervised training. With only a small set of labeled loci images, we can expand surveyed datasets by retrieving semantically similar samples from unexplored imagery. For example, Figure 4 shows a query image with modern buildings and active corrals retrieving samples with comparable features.
Few-shot Locus Segmentation. As shown in Figure 4, we also evaluate semantic segmentation as a downstream task to recognize dense features. The goal is to delineate archaeologically relevant features within an image crop—such as corrals and buildings—while ignoring background elements. This provides pixel-level localization for analyzing settlement patterns and land use. Given the scarcity of labeled archaeological data, we adopt few-shot settings (e.g., 10-30 training samples) to evaluate whether the pre-trained model can adapt to segmentation with minimal annotation. Specifically, we perform few-shot semantic segmentation tasks across three challenging loci: active buildings, active corrals, and archaeological corrals. Similar to [38], patch-level features are extracted from the images via patch tokens and then concatenated with a simple linear segmentation head to generate output logits.